LHC 2011 - latest news

Week 9: Machine coordinators: Ralph Assmann – Bernhard Holzer (until Friday 4/3)

Nine o'clock meetings on the weekend at nine o'clock.

LHC coordination vistar

Commissioning outline



Link to report from morning meeting

Monday 7th March

  • 00h09: Start of transverse damper studies.
  • 04h00: Loss maps.
  • 06h00: Ramp without tune feedback but with radial mod.
  • 08h00: Start injection and dump setup at 450 GeV.

8:30 meeting

14:00 LPC report

Sunday 6th March

  • 00h30: Loss maps.
  • 05h30: Loss maps done. Prepare beam for collimation setup at 3.5 TeV.
  • 09h02: Ramp started. 1 nominal bunch per beam.
  • 10h40: Start collimation setup at 3.5 TeV.
  • 16h04: Dump b1. LBDS self trigger.
  • 17h56: Beam 2 dumped: losses from collimation setup.
  • 18h51: Access for beam dump.
  • 21h07: Injection.
  • 21h53: Machine protection and other tests.

9:00 meeting

Saturday 5th March

  • 00h46: Start ramp to 3.5 TeV. Squeeze to 3.5 m.
  • 03h06: Lost beams.
  • 04h01: Injection.
  • 04h41: Dump problem solved.
  • 05h09: Controller problem DQAMG_4.RQTL11.R2B1.
  • 06h36: QPS problem solved. Injection. Ramp. Squeeze. Beta beat and coupling correction ongoing.
  • 09h00 meeting: OK to go ahead with nominal bunch intensity from beam dump.
  • 10h20: Beam dump with RF interlock from frequency trim.
  • 11h20: Injection nom. bunch. Ramp.
  • 12h53: At 3.5 TeV with 1 nominal bunch per beam.
  • 14h53: Dump. Ramp-down and recover.
  • 15h09: RF module 2B2 down. RF piquet.
  • 16h40: Ready for injection.
  • 17h01: Start injection protection setup.

9:00 meeting

Friday 4th March

  • 08h32: Access for beam dump system.
  • 09h10: Access finished. Beam dump tests without and with beam.
  • 16h00: Start of 2-shift “quiet” beam period for observing beam dump stability (pilot bunch only, ramp OK).
  • 16h39: Observe effect of Q feedback on chroma.
  • 16h48: Start ramp to 3.5 TeV. Squeeze steps for beta beat correction.
  • 22h02: Lost both beams. Cryo gone for MSL4. And the undulator. RU.L4 QPS_OK lost.

8:30 meeting

Thursday 3rd March


8:30 meeting

Wednesday 2nd March

  • 04h36: Cryo problem in arc 81. Access.
  • 09h14: Injection.
    TI8 injection collimator setup.
    Testing new sequence.
    New version of vacuum application (S. Blanchard).
    Updates and modifications of orbit and tune feedbacks.
  • 14h34: Access in PT 2 and 8 for power converters. ATLAS access in shadow.
  • 15h58: RB.A81 lost QPS OK when preparing for precycle. Access required. QPS problem due to loss of communication from beam losses at injection. New software.
    - RQTL111.L3B2 and RCBXH3.R1 tripped, idle at standby. Reset.
    - Access for damper problem in the SPS (3h).
    - LHC-b and CMS accesses in shadow.
    - LBDS access in shadow.
  • 18h37: Pre-cycle. Lost power permit for S67, S78 and S81. Likely from switching off access during pre-cycle (known problem?).
  • 19h34: RSD2.A78B1 tripped during pre cycle.
  • 20h24: One Roman Pot not completed out à interlock. Moved out by TOTEM.
  • 20h43: Injection. 2 pilot bunches per beam.
  • 21h34: Incorporation problems before ramp. Reference mis-match.
  • 22h10: Start of ramp.
  • 22h34: Squeeze.

  • Beam time: ~10h
    Access and recovery: 10h37 (until 20h00)

8:30 meeting

Tuesday 1st March

  • 00h01: MPS tests (RD1.LR1 off). Recover and rampdown.
  • 00h51: Injection. Reboot of OFC and OFSU fixed problem.
  • 02h36: Ramp
  • 03h06: Squeeze. Beta beat correction (B1 complete, B2 to be completed). Beam 2 affected by RQS.A78B2 trips. Coupling not corrected (2 beams required).
  • 08h12: Beam dump and recover.
  • 10h30: Collimation setup.
  • 12h48: Injection protection setup.
  • 13h42: Lost D2L5. Access.
  • 17h43: Injection mode.
  • 19h12: Cryo problem PM25. Access.
  • 22h20: Injection. TL optics measurement and correction.

    Beam time: ~13h
    Access and recovery: 7h09

8:30 meeting

Monday 28th February

  • 8h00: Injection
  • 8h41: Injection and dump protection setup
  • 14h21: XPOC problem
  • 15h28: RF calibration
  • 17h42: Access for RF
  • 20h16: Pre-cycle
  • 21h30: Injection
  • 22h34: Ramp
  • 22h52: Squeeze for orbit checks

8:30 meeting